Can someone help me on this problem?
The length of a rectangular playing field is 5 ft. less than twice its width. If the perimeter of the playing field is 230 ft., I nned to find the length and width of the field.
Thanks,
Sher

Found 2 solutions by stanbon, jenrobrody:
Answer by stanbon(75887)

The length of a rectangular playing field is 5 ft. less than twice its width. If the perimeter of the playing field is 230 ft., I nned to find the length and width of the field.
--------------
Let the width be "x" ft.
Then the length is "2x-5"
Perimeter = 2(length+width)
230 = 2(2x-5+x)
230=2(3x-10)
115=3x-10
3x=125
x=41 2/3 ft (This is the width)
2x-5= 78 1/3 ft (This is the length)
Cheers,
Stan H.

Answer by jenrobrody(19)

let w=width
and since length is 5 less than twice the width,
length=2 times the width minus 5 or
length=2w-5
Perimeter=2*length+2*width:
Perimeter=2(2w-5)+2w=230
distribute:4w-10+2w=230
combine like terms:6w-10=230
add 10 to both sides: 6w=240
divide by 6: w=40
so length= 2(40) - 5 = 80 - 5 = 75
check: 2*(40) + 2*(75) = 80 +150 =230